Review VPartA A parallel-plate capacitor is made from two aluminum-foil sheets, each 6.3 cm wide and 6.0 m long. Between the sheets is a Teflon strip of the same width and length that is 4.5x10-2 mm thick. What is the capacitance of this capacitor? (The dielectric constant of Teflon is 2.1.) Express your answer using two significant figures. Submit Request Answer
Answer #1

Area

A=0.063*6 =0.378 m2

Capacitance of the capacitor is given by

2.1 × 8.8542 × 10-12 × 0.378 (4.5 × 10-2 × 10-3) KeoA

C=1.56*10-7 F =0.156 uF

C=0.16 uF (approx up to 2 significant figures)